# EXTENSIONS OF UNIRATIONAL GROUPS

### v2

@@ 1 Introduction
In [Ach1, Question 4.9], Achet posed the following question: Is a commutative extension of two unirational algebraic $K$-groups necessarily also unirational? His motivation lies in the study of the maximal unirational $K$-subgroup $G_{\mathrm{uni}}$ of an algebraic group, as an affirmative answer implies that, for commutative $G$, the group $G/G_{\mathrm{uni}}$ admits no nontrivial unirational subgroup. The question trivially admits an affirmative answer when $K$ is perfect, as in this case all connected linear algebraic $K$-groups are unirational [Bor, Ch. V,Th. 18.2(ii)].
@@
In this paper, we answer Achet’s question, and more generally undertake a study of extensions of unirational algebraic groups. In §2, we prove that every extension of unirational groups is unirational over fields of degree of imperfection $1$ (Theorem 2.4). However, this fails over every field of degree of imperfection >1>1. In fact, in §3 we construct over every such field a commutative $p$-torsion extension of unirational wound unipotent groups which is not unirational (Example 3.4). We also construct a non-unirational commutative $p$-torsion wound unipotent extension of the additive group by a unirational wound unipotent group (Example 3.5). Thus Achet’s question has a negative answer over every field of degree of imperfection >1>1.
@@
Finally, in §5, we continue with the theme of deducing unirationality properties from quotients, but for unirationality itself rather than ext-unirationality. More precisely, we prove that over a field $K$ of degree of imperfection $r$, a smooth connected unipotent $K$-group $U$ is unirational if and only if the quotient $U/\mathscr{D}_{r}U$ of $U$ by the $r$th term in its lower central series is (Theorem 5.7). Of particular note is the case $r\leq 2$, which tells us that, when $K$ has degree of imperfection $\leq 2$, unirationality of $U$ is equivalent to that of its maximal commutative quotient $U^{\operatorname{\rm{ab}}}$. We show that this is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent $K$-group $U$ such that $U^{\operatorname{\rm{ab}}}$ is unirational, but $U$ is not (Example 5.9).
@@
Throughout this paper, $K$ denotes a field, and when it appears, $p$ denotes a prime number equal to the characteristic of $K$. By a linear algebraic $K$-group, or a linear algebraic group over $K$, we shall mean a smooth affine $K$-group scheme. Recall that the degree of imperfection of a field $K$ of characteristic $p$ is defined to be $\mathrm{log}_{p}[K:K^{p}]$. This quantity is always either infinite or a nonnegative integer. The degree of imperfection of $K$ also equals $\mathrm{dim}_{K}(\Omega^{1}_{K/\mathbf{F}_{p}})$ as well as the maximal size of a $p$-independent subset of $K$ [Mat, Th. 26.5].
@@ 2 Degree of imperfection 11
In this section we will prove that an arbitrary (not necessarily commutative) extension of unirational algebraic $K$-groups must be unirational when $K$ has degree of imperfection $1$. We begin by tying up a loose end concerning a basic permanence property of permawound unipotent groups. Recall that a smooth unipotent $K$-group scheme $U$ is called permawound when, for any exact sequence of finite type $K$-group schemes
@@
$E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$.
@@
If $K$ is perfect, then the assertion is trivially true by [Ros2, Prop. 5.2(i)], so assume that $K$ is imperfect. Then $U^{\prime}$ and $U^{\prime\prime}$ are connected [Ros2, Prop. 6.2], hence so is $U$. We proceed by dimension induction. Note first that, since every quotient of $U$ is an extension of a quotient of $U^{\prime\prime}$ by a quotient of $U^{\prime}$, and since permawoundness is (by definition) inherited by quotients, it follows that we may assume that every quotient of $U$ by a positive-dimensional normal $K$-subgroup is permawound. Consider first the case in which $U$ is not wound. Suppose that we have an exact sequence of finite type $K$-groups
@@
Let $\overline{U}:=U/V$, let $\overline{U}_{s}\subset\overline{U}$ denote the maximal split subgroup, and let $\overline{U}_{w}:=\overline{U}/\overline{U}_{s}$ denote the wound quotient, which is permawound because $V$ is nontrivial. Because $V\subset U$ is central, the commutator map $U\times U\rightarrow U$ descends to a map $c\colon\overline{U}\times\overline{U}\rightarrow U$. We claim that $c$ further descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$. Indeed, let $u_{1},u_{2}\in\overline{U}(K_{s})$. Then $c$ restricts to a map $u_{1}\overline{U}_{s}\times u_{2}\overline{U}_{s}\rightarrow U_{K_{s}}$ which must be constant because $\overline{U}_{s}^{2}$ is split while $U_{K_{s}}$ is wound. Because $\overline{U}^{2}(K_{s})$ is Zariski dense in $\overline{U}^{2}$, it follows that $c$ is $\overline{U}_{s}$-invariant in both arguments, hence $c$ indeed descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$ which we by abuse of notation also denote by $c$. Because $\overline{U}_{w}$ is wound and permawound, it is commutative [Ros3, Cor. 10.4], so $c$ is the constant map to $1\in U(K)$ by [Ros3, Lem. 10.2]. It follows that $U$ is commutative. Because commutative permawound groups are trivially weakly permawound [Ros2, Def. 5.1], $U^{\prime}$ and $U^{\prime\prime}$ are weakly permawound, hence so is $U$ [Ros2, Prop. 5.6]. It follows that $U$ is permawound [Ros2, Prop. 6.9]. ∎
@@
For the if direction, use [Ros3, Th. 1.9(i)]. For the converse, [Ros3, Prop. 9.6] yields the result when $U$ is wound. For arbitrary $U$, unirationality implies that $U$ is smooth and connected. Then we have an exact sequence
@@
with $U_{s}$ split and $U_{w}$ wound. Because $U$ is unirational, so is $U_{w}$, which is therefore permawound. Then [Ros2, Prop. 5.3] and Proposition 2.1 show that $U$ is permawound. ∎
@@
We may assume that $K$ is separably closed [Ros3, Th. 1.6]. Let $f\colon X\rightarrow U$ be a dominant map from a dense open subscheme $X\subset\mathbf{A}^{n}$. Then $G$ is in particular a $T$-torsor over $U$, and $f^{*}(G)$ is a $T$-torsor over $X$, which is trivial because $T$ is split and $\operatorname{Pic}(X)=0$. Thus $f^{*}(G)$ is rational, so the dominant map $f^{*}(G)\rightarrow G$ proves that $G$ is unirational. ∎
@@
with exactness on the right using [Bor, Ch. IV,11.14,Cor. 2]. The groups $Z_{G^{\prime\prime}}(T^{\prime\prime})$ and $Z_{G^{\prime}}(T^{\prime})$ are unirational [Ros3, Prop 7.12], hence so are the two groups on the end in (2.1). They are also unipotent, the unipotence a consequence of the maximality of the respective tori combined with [Bor, Ch. IV,11.5,Cor. 2]. Now an extension of unirational unipotent groups over a field of degree of imperfection $1$ is still unirational by combining Propositions 2.2 and 2.1. Thus $Z_{G}(T)/T$ is unirational, hence so is $Z_{G}(T)$ (Lemma 2.3), hence so is $G$ [Ros3, Prop. 7.12]. ∎
@@ 3 Examples of extensions of unirational groups that are not unirational
In this section we will show by example – over every field of degree of imperfection >1>1 – that unirationality is not inherited by extensions in general. Our examples will be of two main types: Our first example will be an extension of a wound unirational group by a wound unirational group, while our second will be an extension of ${\mathbf{G}}_{\rm{a}}$ by a wound unirational group. Of course, any extension of a unirational group by a split unipotent one is unirational, as ${\rm{H}}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$. Our examples will all be commutative and $p$-torsion, hence they provide a negative answer to Achet’s question over fields of degree of imperfection >1>1. The proofs that these examples work will require various auxiliary lemmas, to which we now turn.
@@
Note in particular that $V_{n,\alpha}$ is unirational. This also may be seen as folllows: $V_{n,\alpha}$ is defined over the field $\mathbf{F}_{p}(\alpha)$, so we may verify its unirationality over this field, which has degree of imperfection $1$. But over $\mathbf{F}_{p}(\alpha)$, $V_{n,\alpha}$ is permawound [Ros2, Props. 6.4,6.9,3.5], hence unirational [Ros3, Th. 1.9(i)].
@@
By [Oes, Ch. VI,Prop. 5.3], this last equation describes the $K_{n-1}$-group $\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}$, so via (3.2), we obtain a nonzero $K_{n-1}$-homomorphism $V_{K_{n-1}}\rightarrow\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}$, whence a nonzero homomorphism $g\colon V\rightarrow\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ over $K$. It remains to prove that this (and indeed, any such) homomorphism is an isomorphism. The principal part of the polynomial (3.1) defining $V$ is reduced and universal over $F$ (for instance, using [Ros2, Prop. 3.5(ii)]), so $V$ is permawound over $F$ [Ros2, Props. 6.4,6.9], hence unirational [Ros3, Th. 1.9(i)]. Thus $\operatorname{im}(g)\subset\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is a nonzero unirational $K$-subgroup. But the latter group is a minimal unirational $K$-group in the sense that its only nonzero unirational $K$-subgroup is itself [Ach2, Props. 2.18,2.20], hence $g$ is surjective. To prove that $g$ is injective, we note that dimension considerations imply that $\ker(g)$ is finite. Because $V/\ker(g)\simeq\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is wound unipotent, $\ker(g)$ is weakly permawound [Ros2, Prop. 5.5], hence has strictly positive dimension if nontrivial (for instance, by [Ros2, Th. 9.5]; recall that weak permawoundness is insensitive to algebraic separable extension of the ground field [Ros2, Prop. 6.7]). Thus $\ker(g)=0$, so $g$ is an isomorphism. ∎
@@
Because $U$ is unirational, it is generated by maps from nonempty open subschemes of $\mathbf{P}^{1}$ with a rational point lying above the identity. But any such map $X\rightarrow U$ factors through a homomorphism $\operatorname{R}_{D/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ for some finite closed subscheme $D\subset\mathbf{P}^{1}$ by [BLR, §10.3, Thm. 2]. (For a more detailed explanation of how this follows from that theorem, see the discussion before Theorem 2.2 in [Ros1, §2].) Because $U$ is wound, $D$ may be taken to be reduced by [Ros1, Lem. 2.3]. Write $D=\coprod\{x_{i}\}$ with the $x_{i}$ (reduced) closed points in $\mathbf{P}^{1}$. Then $\operatorname{R}_{D/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{i}\operatorname{R}_{K(x_{i})/K}({\mathbf{G}}_{\rm{m}})$. Hence $U$ is generated by maps from $\operatorname{R}_{K(x)/K}({\mathbf{G}}_{\rm{m}})$ as $x$ varies over closed points of $\mathbf{P}^{1}$. Because $K$ is separably closed, $K(x)/K$ is a (finite) purely inseparable extension, and it is a primitive extension (generated by a single element) by [Ros3, Lem. 5.4]. We therefore conclude that $U$ is generated by maps from $\operatorname{R}_{L/K}({\mathbf{G}}_{\rm{m}})$ as $L$ varies over primitive nontrivial purely inseparable extensions of $K$. But any such extension is of the form $K(\alpha^{1/p^{n}})$ for some $\alpha\in K-K^{p}$ and $n\geq 1$. The multiplication by $p$ map on $\operatorname{R}_{K(\alpha^{1/p^{n}})/K}({\mathbf{G}}_{\rm{m}})$ has image $\operatorname{R}_{K(\alpha^{1/p^{n-1}})/K}({\mathbf{G}}_{\rm{m}})$, as may, for instance, be checked on $\mathbf{F}_{p}(\alpha)_{s}$-points, since both groups are smooth and defined over $\mathbf{F}_{p}(\alpha)$. Thus, because $U$ is $p$-torsion, it is generated by maps from groups of the form
@@
which is isomorphic to $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ [Ros2, Prop. 7.4].
By [Ros2, Lem. 7.9], therefore, and an easy induction, any map $V_{n,\alpha}\rightarrow V_{1,\lambda}$ must factor through the map $V_{n,\alpha}\rightarrow V_{1,\alpha}$ obtained by iterating the maps $f_{m}$ above. The lemma therefore reduces to the case $n=1$, where we must show that the map in question is obtained by a linear change of variables, and that the coefficient of $S$ in $X$ is nonzero. (That $X$ may be written uniquely as a $p$-polynomial of degree $\leq 1$ in $S$ follows from [Ros1, Prop. 6.4].) For this, we may extend scalars to $K_{s}$ and thereby assume that $K$ is separably closed. The lemma then follows from [Ros2, Lem. 7.1,Prop. 9.7], since, for a linear change of variables that transforms the polynomial $X+\lambda^{p-1}X+\sum_{i=0}^{p-2}\lambda^{i}S_{i}^{p}$ into a nonzero scalar multiple of $S+\alpha^{p-1}S+\sum_{i=0}^{p-2}\alpha^{i}S_{i}^{p}$, $X=X(S,S_{i})$ must be a nonzero multiple of $S$. ∎
@@
Applying [Ros1, Prop. 6.4], there exist $p$-polynomials $X_{0},\dots,X_{p-1}\in K[S,(S_{j})_{j}]$ of degree $\leq 1$ in $S$ such that
@@
as homomorphisms $V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}$. Applying [Ros1, Prop. 6.4] once more, if we use the equation (3.1) to eliminate $S^{p}$ from the $X_{i}^{p}$ terms above, then this becomes an identity of polynomials in $S$ and the $S_{j}$. Write

### v3

@@ 1 Introduction
Recall that a finite type $K$-scheme $X$ is said to be unirational when there is a dominant rational map $\mathbf{A}^{n}\rightarrow X$ for some $n>0$. By dominant, we mean in the sense that the schematic closure is all of $X$, so in particular $X$ must be geometrically integral. Unirational varieties are of interest for a variety of reasons, one of which is that they admit many (a Zariski dense set of) rational points, at least when $K$ is infinite. Unirational algebraic groups, in particular, have received attention from various authors; see, for instance, [BLR, §10.3] (on the relationship between unirationality and the existence of Néron models for algebraic groups), [CGP, §11.3] (largely on the relationship between unirationality and various arithmetic properties of algebraic groups), [Ach2] (regarding geometric properties of unirational groups), [Ros1, §6] (about the Picard groups of unirational groups), and [Ros4] (concerned primarily with various rigidity properties exhibited by unirational groups, in some sense analogous to the well-known rigidity properties of abelian varieties). Over perfect fields, every connected linear algebraic groups is unirational [Bor, Ch. V,Th. 18.2(ii)], but this is false over every imperfect field [CGP, Ex. 11.3.1], and so unirationality becomes a nontrivial condition on (connected) linear algebraic groups in the imperfect setting. If $G$ is a connected linear algebraic group, and $T\subset G$ is a maximal $K$-torus, then $G$ is unirational precisely when the unipotent group $Z_{G}(T)/T$ is, where $Z_{G}(T)$ denotes the centralizer of $T$. (Combine [Ros4, Prop. 7.12] and Lemma 2.3 below.) For this reason (and others), unirationality questions for algebraic groups often reduce to the unipotent setting, and thus the theory of unipotent groups plays a central role in such problems.
@@
In [Ach1, Question 4.9], Achet posed the following question: Is a commutative extension of two unirational algebraic $K$-groups necessarily also unirational? His motivation lies in the study of the maximal unirational $K$-subgroup $G_{\mathrm{uni}}$ of an algebraic group. (This subgroup of $G$ may also be described as the group generated by all maps from unirational varieties into $G$ which pass through the identity.) Achet posed his question because an affirmative answer implies that, for commutative $G$, the group $G/G_{\mathrm{uni}}$ admits no nontrivial unirational subgroup. The question trivially admits an affirmative answer when $K$ is perfect, since all connected linear algebraic groups are unirational in this case, but over imperfect fields it becomes highly nontrivial.
In this paper, we answer Achet’s question, and more generally undertake a study of extensions of unirational algebraic groups. In §2, we prove that every extension of unirational groups is unirational over fields of degree of imperfection $1$ (Theorem 2.4). This makes crucial use of the concept of permawound unipotent groups introduced in [Ros3]. We will remind the reader of the definition of permawoundness in §2, but their utility arises not from their definition, but from the fact that they simultaneously exhibit two important properties: One the one hand, they enjoy a certain universality or ubiquity – every commutative $p$-torsion wound unipotent group may be embedded into a permawound one – so that many questions about arbitrary unipotent groups reduce to the permawound case. On the other hand, they exhibit an extraordinary rigidity: After passage to a suitable finite separable extension, every permawound group admits a filtration with successive quotients all isomorphic to one of two particular groups, very much analogously to the manner in which split unipotent groups always (by definition) admit a filtration with ${\mathbf{G}}_{\rm{a}}$ quotients [Ros3, Th. 1.4,1.5]. One sign that they are relevant for our investigations here is that such groups are always unirational [Ros4, Th. 1.9(i)]. In fact, we shall show that a smooth unipotent group over a field of degree of imperfection $1$ is unirational precisely when it is permawound (Proposition 2.2), from which the stability of unirationality under extensions will follow easily.
@@
However, this stability fails over every field of degree of imperfection >1>1. In fact, in §3 we construct over every such field a commutative $p$-torsion extension of unirational wound unipotent groups which is not unirational (Example 3.4). We also construct a non-unirational commutative $p$-torsion wound unipotent extension of the additive group by a unirational wound unipotent group (Example 3.5). Thus Achet’s question has a negative answer over every field of degree of imperfection >1>1.
@@
Finally, in §5, we continue with the theme of deducing unirationality properties from quotients, but for unirationality itself rather than ext-unirationality. More precisely, we prove that over a field $K$ of degree of imperfection $r>1$, a smooth connected unipotent $K$-group $U$ is unirational if and only if the quotient $U/\mathscr{D}_{r}U$ of $U$ by the $r$th term in its lower central series is (Theorem 5.7). Of particular note is the case $r\leq 2$, which – in conjunction with Corollary 4.5 and the unirationality of all connected linear algebraic groups when $r=0$ – tells us that, when $K$ has degree of imperfection $\leq 2$, unirationality of $U$ is equivalent to that of its maximal commutative quotient $U^{\operatorname{\rm{ab}}}$. We show that this is optimal by giving, over every field of degree of imperfection $\geq 3$, an example of a wound unipotent $K$-group $U$ such that $U^{\operatorname{\rm{ab}}}$ is $p$-torsion and unirational even though $U$ fails to be unirational. (Example 5.9).
@@
Throughout this paper, $K$ denotes a field, and when it appears, $p$ denotes a prime number equal to the characteristic of $K$. By a linear algebraic $K$-group, or a linear algebraic group over $K$, we shall mean a smooth affine $K$-group scheme. Recall that the degree of imperfection of a field $K$ of characteristic $p$ is defined to be $\mathrm{log}_{p}[K:K^{p}]$. This quantity is always either infinite or a nonnegative integer. The degree of imperfection of $K$ also equals $\mathrm{dim}_{K}(\Omega^{1}_{K/\mathbf{F}_{p}})$, as well as (tautologically, from the first definition) the size of any $p$-basis of $K$ [Mat, Th. 26.5].
Acknowledgments
I thank Anis Zidani for pointing out an error in an earlier version of Theorem 5.7.
@@ 2 Degree of imperfection 11
In this section, we will prove that an arbitrary (not necessarily commutative) extension of unirational algebraic $K$-groups must be unirational when $K$ has degree of imperfection $1$. We begin by tying up a loose end concerning a basic permanence property of permawound unipotent groups, which were introduced in [Ros3]. Recall [Ros3, Def. 1.2] that a smooth unipotent $K$-group scheme $U$ is called permawound when, for any exact sequence of finite type $K$-group schemes
@@
$E$ contains a copy of ${\mathbf{G}}_{\rm{a}}$; that is, there is an injective (or equivalently, there is a nonzero) $K$-homomorphism ${\mathbf{G}}_{\rm{a}}\rightarrow E$. As we shall show below (Proposition 2.2), it is an easy consequence of the results of [Ros3] and [Ros4] that permawoundness is equivalent to unirationality for unipotent groups over fields of degree of imperfection $1$. At least for unipotent groups, then, we will deduce the permanence of unirationality under extensions from the corresponding property of permawoundness, to be proved presently. The case of general algebraic groups will then be obtained by reduction to the unipotent case, using the fact, a consequence of [Ros4, Prop. 7.12] and Lemma 2.3 below, that $G$ is unirational if and only if the same holds for $Z_{G}(T)/T$, where $T\subset G$ is a maximal torus.
@@
If $K$ is perfect, then the assertion is trivially true by [Ros3, Prop. 5.2(i)], so assume that $K$ is imperfect. Then $U^{\prime}$ and $U^{\prime\prime}$ are connected [Ros3, Prop. 6.2], hence so is $U$. We proceed by dimension induction. Note first that, since every quotient of $U$ is an extension of a quotient of $U^{\prime\prime}$ by a quotient of $U^{\prime}$, and since permawoundness is (by definition) inherited by quotients, it follows that we may assume that every quotient of $U$ by a positive-dimensional normal $K$-subgroup is permawound. Consider first the case in which $U$ is not wound. Suppose that we have an exact sequence of finite type $K$-groups
@@
Let $\overline{U}:=U/V$, let $\overline{U}_{s}\subset\overline{U}$ denote the maximal split subgroup, and let $\overline{U}_{w}:=\overline{U}/\overline{U}_{s}$ denote the wound quotient, which is permawound because $V$ is nontrivial. Because $V\subset U$ is central, the commutator map $U\times U\rightarrow U$ descends to a map $c\colon\overline{U}\times\overline{U}\rightarrow U$. We claim that $c$ further descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$. Indeed, let $u_{1},u_{2}\in\overline{U}(K_{s})$. Then $c$ restricts to a map $u_{1}\overline{U}_{s}\times u_{2}\overline{U}_{s}\rightarrow U_{K_{s}}$ which must be constant because $\overline{U}_{s}^{2}$ is split while $U_{K_{s}}$ is wound. Because $\overline{U}^{2}(K_{s})$ is Zariski dense in $\overline{U}^{2}$, it follows that $c$ is $\overline{U}_{s}$-invariant in both arguments, hence $c$ indeed descends to a map $\overline{U}_{w}\times\overline{U}_{w}\rightarrow U$ which we by abuse of notation also denote by $c$. Because $\overline{U}_{w}$ is wound and permawound, it is commutative [Ros4, Cor. 10.4], so $c$ is the constant map to $1\in U(K)$ by [Ros4, Lem. 10.2]. It follows that $U$ is commutative. Because commutative permawound groups are trivially weakly permawound [Ros3, Def. 5.1], $U^{\prime}$ and $U^{\prime\prime}$ are weakly permawound, hence so is $U$ [Ros3, Prop. 5.6]. It follows that $U$ is permawound [Ros3, Prop. 6.9]. ∎
@@
For the if direction, use [Ros4, Th. 1.9(i)]. For the converse, [Ros4, Prop. 9.6] yields the result when $U$ is wound. For arbitrary $U$, unirationality implies that $U$ is smooth and connected. Then we have an exact sequence
@@
with $U_{s}$ split and $U_{w}$ wound. Because $U$ is unirational, so is $U_{w}$, which is therefore permawound. Then [Ros3, Prop. 5.3] and Proposition 2.1 show that $U$ is permawound. ∎
@@
We may assume that $K$ is separably closed [Ros4, Th. 1.6]. Let $f\colon X\rightarrow U$ be a dominant map from a dense open subscheme $X\subset\mathbf{A}^{n}$. Then $G$ is in particular a $T$-torsor over $U$, and $f^{*}(G)$ is a $T$-torsor over $X$, which is trivial because $T$ is split and $\operatorname{Pic}(X)=0$. Thus $f^{*}(G)$ is rational, so the dominant map $f^{*}(G)\rightarrow G$ proves that $G$ is unirational. ∎
We come now to the main result of this section.
@@
with exactness on the right using [Bor, Ch. IV,11.14,Cor. 2]. The groups $Z_{G^{\prime\prime}}(T^{\prime\prime})$ and $Z_{G^{\prime}}(T^{\prime})$ are unirational [Ros4, Prop 7.12], hence so are the two groups on the end in (2.1). They are also unipotent, the unipotence a consequence of the maximality of the respective tori combined with [Bor, Ch. IV,11.5,Cor. 2]. Now an extension of unirational unipotent groups over a field of degree of imperfection $1$ is still unirational by combining Propositions 2.2 and 2.1. Thus $Z_{G}(T)/T$ is unirational, hence so is $Z_{G}(T)$ (Lemma 2.3), hence so is $G$ [Ros4, Prop. 7.12]. ∎
@@ 3 Examples of extensions of unirational groups that are not unirational
In this section we will show by example – over every field of degree of imperfection >1>1 – that unirationality is not inherited by extensions in general. Our examples will be of two types: Our first example will be an extension of a wound unirational group by a wound unirational group, while our second will be an extension of ${\mathbf{G}}_{\rm{a}}$ by a wound unirational group. Of course, any extension of a unirational group by a split unipotent one is unirational, as ${\rm{H}}^{1}(X,{\mathbf{G}}_{\rm{a}})=0$ for every affine scheme $X$. Our examples will all be commutative and $p$-torsion, hence they provide a negative answer to Achet’s question over fields of degree of imperfection >1>1. The proofs that these examples work will require various auxiliary lemmas, to which we now turn.
@@
Note in particular that $V_{n,\alpha}$ is unirational. This also may be seen as folllows: $V_{n,\alpha}$ is defined over the field $\mathbf{F}_{p}(\alpha)$, so we may verify its unirationality over this field, which has degree of imperfection $1$. But over $\mathbf{F}_{p}(\alpha)$, $V_{n,\alpha}$ is permawound [Ros3, Props. 6.4,6.9,3.5], hence unirational [Ros4, Th. 1.9(i)].
@@
By [Oes, Ch. VI,Prop. 5.3], this last equation describes the $K_{n-1}$-group $\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}$, so via (3.2), we obtain a nonzero $K_{n-1}$-homomorphism $V_{K_{n-1}}\rightarrow\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}}$, whence a nonzero homomorphism $g\colon V\rightarrow\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ over $K$. It remains to prove that this (and indeed, any such) homomorphism is an isomorphism. The principal part of the polynomial (3.1) defining $V$ is reduced and universal over $F$ (for instance, using [Ros3, Prop. 3.5(ii)]), so $V$ is permawound over $F$ [Ros3, Props. 6.4,6.9], hence unirational [Ros4, Th. 1.9(i)]. Thus $\operatorname{im}(g)\subset\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is a nonzero unirational $K$-subgroup. But the latter group is a minimal unirational $K$-group in the sense that its only nonzero unirational $K$-subgroup is itself [Ach2, Props. 2.18,2.20], hence $g$ is surjective. To prove that $g$ is injective, we note that dimension considerations imply that $\ker(g)$ is finite. Because $V/\ker(g)\simeq\operatorname{R}_{K_{n-1}/K}(\operatorname{R}_{K_{n}/K_{n-1}}({\mathbf{G}}_{\rm{m}})/{\mathbf{G}}_{\rm{m}})$ is wound unipotent, $\ker(g)$ is weakly permawound [Ros3, Prop. 5.5], hence has strictly positive dimension if nontrivial (for instance, by [Ros3, Th. 9.5]; recall that weak permawoundness is insensitive to algebraic separable extension of the ground field [Ros3, Prop. 6.7]). Thus $\ker(g)=0$, so $g$ is an isomorphism. ∎
@@
Because $U$ is unirational, it is generated by maps from nonempty open subschemes of $\mathbf{P}^{1}$ with a rational point lying above the identity. But any such map $X\rightarrow U$ factors through a homomorphism $\operatorname{R}_{D/K}({\mathbf{G}}_{\rm{m}})\rightarrow U$ for some finite closed subscheme $D\subset\mathbf{P}^{1}$ by [BLR, §10.3, Thm. 2]. (For a more detailed explanation of how this follows from that theorem, see the discussion before Theorem 2.2 in [Ros2, §2].) Because $U$ is wound, $D$ may be taken to be reduced by [Ros2, Lem. 2.3]. Write $D=\coprod\{x_{i}\}$ with the $x_{i}$ (reduced) closed points in $\mathbf{P}^{1}$. Then $\operatorname{R}_{D/K}({\mathbf{G}}_{\rm{m}})\simeq\prod_{i}\operatorname{R}_{K(x_{i})/K}({\mathbf{G}}_{\rm{m}})$. Hence $U$ is generated by maps from $\operatorname{R}_{K(x)/K}({\mathbf{G}}_{\rm{m}})$ as $x$ varies over closed points of $\mathbf{P}^{1}$. Because $K$ is separably closed, $K(x)/K$ is a (finite) purely inseparable extension, and it is a primitive extension (generated by a single element) by [Ros4, Lem. 5.4]. We therefore conclude that $U$ is generated by maps from $\operatorname{R}_{L/K}({\mathbf{G}}_{\rm{m}})$ as $L$ varies over primitive nontrivial purely inseparable extensions of $K$. But any such extension is of the form $K(\alpha^{1/p^{n}})$ for some $\alpha\in K-K^{p}$ and $n\geq 1$. The multiplication by $p$ map on $\operatorname{R}_{K(\alpha^{1/p^{n}})/K}({\mathbf{G}}_{\rm{m}})$ has image $\operatorname{R}_{K(\alpha^{1/p^{n-1}})/K}({\mathbf{G}}_{\rm{m}})$, as may, for instance, be checked on $\mathbf{F}_{p}(\alpha)_{s}$-points, since both groups are smooth and defined over $\mathbf{F}_{p}(\alpha)$. Thus, because $U$ is $p$-torsion, it is generated by maps from groups of the form
@@
which is isomorphic to $\operatorname{R}_{K^{1/p}/K}(\alpha_{p})$ [Ros3, Prop. 7.4].
By [Ros3, Lem. 7.9], therefore, and an easy induction, any map $V_{n,\alpha}\rightarrow V_{1,\lambda}$ must factor through the map $V_{n,\alpha}\rightarrow V_{1,\alpha}$ obtained by iterating the maps $f_{m}$ above. The lemma therefore reduces to the case $n=1$, where we must show that the map in question is obtained by a linear change of variables, and that the coefficient of $S$ in $X$ is nonzero. (That $X$ may be written uniquely as a $p$-polynomial of degree $\leq 1$ in $S$ follows from [Ros2, Prop. 6.4].) For this, we may extend scalars to $K_{s}$ and thereby assume that $K$ is separably closed. The lemma then follows from [Ros3, Lem. 7.1,Prop. 9.7], since, for a linear change of variables that transforms the polynomial $X+\lambda^{p-1}X+\sum_{i=0}^{p-2}\lambda^{i}S_{i}^{p}$ into a nonzero scalar multiple of $S+\alpha^{p-1}S+\sum_{i=0}^{p-2}\alpha^{i}S_{i}^{p}$, $X=X(S,S_{i})$ must be a nonzero multiple of $S$. ∎
@@
Applying [Ros2, Prop. 6.4], there exist $p$-polynomials $X_{0},\dots,X_{p-1}\in K[S,(S_{j})_{j}]$ of degree $\leq 1$ in $S$ such that
@@
as homomorphisms $V_{n,\alpha}\rightarrow{\mathbf{G}}_{\rm{a}}$. Applying [Ros2, Prop. 6.4] once more, if we use the equation (3.1) to eliminate $S^{p}$ from the $X_{i}^{p}$ terms above, then this becomes an identity of polynomials in $S$ and the $S_{j}$. Write